Find two consecutive odd integers whose sum is 28
Found 2 solutions by Alan3354, Earlsdon:
Answer by Alan3354(69443)

28/2 = 14, the center
--> 13 & 15

Answer by Earlsdon(6294)

Let x = the first odd integer, then x+2 = the next consecutive odd integer.
x+(x+2) = 28 "...odd integers whose sum is 28." Simplify this to...
2x+2 = 28 Subtract 2 from both sides.
2x = 26 Divide both sides by 2.
x = 13 and...
x+2 = 15
The two integers are 13 and 15